OpenStudy (anonymous):

I have the answer can someone just check me please. Which of the following is f '(x) for f(x)=2x(x^2+5)

OpenStudy (anonymous):

I got 2x^2+10

OpenStudy (whpalmer4):

What is f(x) after applying the distributive property?

OpenStudy (anonymous):

2x^2+10x

OpenStudy (whpalmer4):

2x*x^2= 2x^2?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

So then you get 4x+10x

OpenStudy (anonymous):

is this right

OpenStudy (whpalmer4):

2x*x^2 = something other than 2x^2!

OpenStudy (whpalmer4):

2x*x^2= 2x*x*x = ?

OpenStudy (anonymous):

so the would the answer be 6x+10x

OpenStudy (whpalmer4):

\[2x(x^2+5) = 2x^3+10x\] \[\frac{d}{dx}[2x^3+10x] =\]?

OpenStudy (anonymous):

ok so I just redid it and was wondering if 6x^2+10 would be correct

OpenStudy (whpalmer4):

Ding ding ding, that is correct.